# D-LiFT: Improving LLM-based Decompiler Backend via Code Quality-driven Fine-tuning

### v1

@@ Abstract
Decompilers, which reconstruct human-readable source code from binary executables, are vital to many security tasks. Yet, despite recent advances, their output often suffers from syntactic and semantic errors and remains difficult to read. Recently, with the advent of large language models (LLMs), researchers began to explore the potential of LLMs to refine decompiler output. Nevertheless, our study of these approaches reveals significant limitations, such as introducing new errors and relying on unreliable accuracy validation.
@@
In this paper, we present D-LiFT, an automated decompiler backend that harnesses and further trains LLMs to improve the quality of decompiled code via reinforcement learning (RL). Unlike prior work that overlooks preserving accuracy, D-LiFT adheres to a key principle for enhancing the quality of decompiled code: preserving accuracy while improving readability. Central to D-LiFT, we propose D-Score, an integrated quality assessment system to score the decompiled code from multiple aspects. In line with our principle, D-Score assigns low scores to any inaccurate output and only awards higher scores for readability to code that passes the accuracy check. Specifically, D-Score first verifies the syntactic and semantic correctness via the compiler and symbolic execution; only if a candidate is deemed accurate, it then evaluates readability using established metrics to compare the LLM output with the original decompiled code. The score will then be fed back to the LLM for fine-tuning. Our implementation, based on Ghidra and a range of LLMs, demonstrates significant improvements for the accurate decompiled code from the coreutils and util-linux projects. Compared to baseline LLMs without D-Score-driven fine-tuning, D-LiFT produces 55.3% more improved decompiled functions, as measured by D-Score. Moreover, when selecting the best output among the baseline model, the fine-tuned model, and the original decompiled code, our fine-tuned model yields the best result for 47.3% of functions, whereas the baseline model does so for only 20.9%.
1 Introduction
@@
Binary decompilation is the process of generating high-level source code, with human-comprehensible variables, control flows, and data structures from a binary program. Decompilation has found extensive security applications, such as software reverse engineering [60], vulnerability discovery [40, 22], program patching and hardening [48], and cyber forensics [9]. However, source code generated by decompilers very often has problems in the aspects of syntactic and semantic correctness [37, 72, 11, 16]: A piece of decompiled code may not compile or preserve the semantics of the original binary. Moreover, despite advances in the past decade [7, 70, 12, 35, 71, 67, 66, 8, 33], even with syntactic and semantic correctness, the decompiled code may still be hard for human users to read. In fact, a recent survey [60] found that roughly 70% of reverse engineers still prefer working with the assembly code, due to the incorrectness (i.e., inaccuracy) and/or low readability of the decompiled source code.
@@
In recent years, with the advent of Large Language Models (LLMs), researchers started to explore the potential of LLMs to generate or improve decompiled code [64, 55, 25, 29, 18, 28, 63, 65]. For example, to generate more readable decompiled output, LLM4Decompile [55] introduces supervised fine-tuning (SFT) for the training, and DeGPT [25] introduces a three-role mechanism to help LLMs inference. These LLM-assisted decompilation improvement tools, such as LLM4Decompile and GhidraMCP [31], have rapidly gained attention within the security community. For instance, LLM4Decompile was ranked in 4th place in GitHub Day Trending.
Unfortunately, our analysis of these state-of-the-art, LLM-based decompiled code improvement methods shows significant limitations. For LLM4Decompile, our study (detailed in Section 6.1.3) reveals that, even when the original decompiled code is error-free, new errors are found in 93.2% of its generated functions. For DeGPT, its semantic correctness check is non-deterministic and does not enhance the LLM’s inherent capability to generate more accurate decompiled code. Moreover, our investigation uncovers a unique challenge, multiple valid ground truth possibilities, in improving LLMs for decompilation, as detailed in Section 3.
@@
To improve the overall quality while preserving accuracy of the decompiled code using LLMs, we propose D-LiFT  [^1] D-LiFT” reflects the decompilation pipeline with a “Decompiler” front-end and an “LLM with Fine Tuning” back-end., an automatic decompilation pipeline with a decompiler front-end and an LLM back-end, which is fine-tuned to improve the quality of the decompiled code. D-LiFT takes a policy model (i.e., baseline model), a decompiler, and a set of real-world binaries as inputs, and provides a code quality-enhanced model to generate more readable decompiled code while preserving accuracy. Unlike prior work that overlooks preserving accuracy, D-LiFT follows a key principle for improving the quality of decompiled code: preserving accuracy while improving readability. To this end, we propose D-Score, a novel code quality assessment function, which provides comprehensive scoring for LLM-generated decompiled code based on accuracy and readability measurements. Specifically, to assess the accuracy of decompiled code, D-Score employs a compiler to generate the syntax feedback and symbolic execution to compare the semantics of the generated code against the corresponding function in the original binary. To assess the readability of decompiled code, D-Score computes a score by applying two established readability metrics to compare the LLM’s output with the original decompiled code. By leveraging D-Score, D-LiFT enhances the LLMs’ ability to generate higher-quality decompiled code.
@@
We implement D-LiFT using Ghidra [3] and a number of LLMs, and evaluate D-LiFT using functions from the coreutils [13] and util-linux [58] projects. Overall, as measured by D-Score, D-LiFT can improve the quality of 55.3% more functions, compared to the baseline LLMs without D-Score-driven fine-tuning. Interestingly, we observe a significant “improve-ability gap” between the accurate and inaccurate decompiled functions originally produced by the decompiler. All LLMs, including baseline and fine-tuned models, face challenges in improving inaccurate decompiled code, with improvement rates of just 8.02% for the originally inaccurate functions compared to 86.2% for the originally accurate ones. Moreover, for functions that are accurately decompiled, when choosing the best output among the baseline model (i.e., decompiler+LLM without fine-tuning), the fine-tuned model (i.e., D-LiFT), and the original decompiled code, we find that on average, 47.3% of the top-scoring functions come from D-LiFT, whereas only 20.9% come from the baseline model. Overall, our main contributions are:
@@
  • We design D-LiFT, an automatic decompilation pipeline with an LLM-based back-end that is fine-tuned using reinforcement learning to improve the quality of the decompiled code, adhering to the principle of preserving accuracy while improving readability.
@@
  • We propose D-Score, an integrated scoring mechanism designed specifically for decompilation recovery tasks. The framework incorporates existing analytical tools and proven metrics to deliver a balanced evaluation of decompiled code, assessing both syntactic and semantic correctness as well as readability properties to provide effective training feedback to the LLM.
@@
  • We implement D-LiFT based on Ghidra and fine-tune three LLMs, and achieve significant decompiled code improvement for widely used benchmark functions.
2 Background and Motivation
2.1 LLM application and training
2.1.1 Applications of LLMs
@@
Large language models (LLMs) have seen wide adoption and investigation, particularly in the domain of code generation. Industry has released numerous LLM-powered tools, e.g., GitHub Copilot [19], Google Gemini Code Assistant [20], and Meta LLaMA Coder [50], to streamline different programming tasks. Academia has mirrored this enthusiasm: the number of publications on code generation rose from 11 in 2022 to 75 in 2023 and then to 140 in 2024—a 1,272% increase [27]. In particular, LLMs have demonstrated promising capabilities across a range of code-to-code generation tasks, including code translation, code completion, automated bug fixing (code repair), and mutant generation. In the security domain, beyond LLM4Decompile, reverse-engineering teams have also leveraged LLMs via Model Context Protocol (MCP) servers to infer variable and function names, producing more readable decompiled output; one such tool, GhidraMCP [31], garnered 4.3k stars on GitHub within its first month after release.
2.1.2 Training LLMs
@@
Instruction tuning runs the supervised learning paradigm aiming to align the model’s output with a single desired completion. For instance, LLM4Decompile applies Supervised Fine-Tuning (SFT) to align the desired output with the original source code.
@@
Reinforcement learning, on the other hand, refines the model through iterative actions and feedback: the model generates candidate output(action), which will be assessed by a reward function that provides feedback, guiding the model toward higher-quality results. In code generation, the feedback usually comes from frameworks, such as unit tests and compilers. Some RL policies, such as PPO [52], also require a single ideal completion to train a critic model that predicts long-term rewards, helping the model choose better actions over time. More recently, Group Relative Policy Optimization (GRPO) [53] achieved promising results on math-related tasks. By normalizing the rewards of candidate outputs instead of relying on a critic model, GRPO eliminates the limitation of accepting only one ideal completion.
2.2 Decompiled Code Accuracy
The accuracy of the decompiled code, including recompilation syntactic failures and semantic deviations between the decompiled code and the original source code, was first systematically evaluated using Equivalence Modulo Input (EMI) tests [37]. Since then, various methods, including symbolic execution [72], random testing [11], fuzzing [68], and manual inspection [16], have been employed to assess decompiler accuracy. Within these studies, the definition of semantic deviation varies across studies. For example, one [72] may compare function return values between the decompiled code and the binary, while the other [11] may instrument global variables and employ checksum comparisons to evaluate semantic correctness against the original source code.
2.3 Decompiled Code Readability
Quantitative assessment of code readability has long been a topic in software engineering. Buse and Weimer (B&W) [10] led the way by recruiting 120 human evaluators to rate 100 short code snippets, then building a mathematical model, featuring metrics like the number of variables per function, to evaluate the readability.
In the decompilation field, R2I [17] was the first approach to measure readability specifically for decompiled code. It defines a set of features that can be categorized into five feature groups: code quality, user preference, conflicting features, erroneous syntax, and general features. By analyzing these features extracted from the code, it provides relative readability scores across different decompilers. Note that, since R2I focuses on the comparison across different decompilers, many of its features cannot be generically applied to the LLM-generated code.
2.4 Motivations
@@
Although LLM-based code generation is increasingly prevalent, using it to improve decompiled code still faces major challenges.
@@
LLM causes inaccuracy of decompiled code. Though LLMs are widely used for code enhancement [27], they are known to introduce inaccuracies in code generation tasks [46]. In the decompiled-code improvement scenario, we also observe that LLM-refined output frequently shows new errors. Here, we follow the previous work to define the inaccuracy in the decompiled code as either syntax errors that prevent successful compilation or semantic deviations between the LLM-generated code and the original binary’s behavior. As shown in Table III, an average 44.2% of functions that were originally accurate become inaccurate after LLM processing. For the root cause, as the example shown later in Figure 8, we observed that LLMs frequently make errors on small details, such as omitting brackets or instructions, or referencing the wrong variable, that are hard to spot yet vital for accuracy.
@@
These observations reveal that though LLMs may improve readability, they often introduce incorrectness of decompiled code, a factor that prior studies [72, 37, 11, 16, 68] have identified as vital for effective decompilation. Hence, we propose an overarching principle for improving decompiled code quality: preserving accuracy while improving readability.
However, none of the existing work improves LLMs’ code correction (i.e., accuracy) capability. Specifically, LLM4Decompile, while effective at boosting readability, compromises the accuracy of the decompiled code, as shown in Table III. DeGPT, meanwhile, is unable to improve the LLM’s inherent generation capabilities, because its fixes occur only at the inference stage rather than during training. These insights lead us to design a framework that enhances LLMs’ ability to correct errors in the decompiled code while improving its readability.
3 Design Challenges
Section  2.4 motivates the need to enhance the LLM in decompiled code improvement. In this section, we summarize the design-level challenges.

### v2

@@ Abstract
As one of the key tools in many security tasks, decompilers reconstruct human-readable source code from binaries. Yet, despite recent advances, their outputs often suffer from syntactic and semantic errors and remain difficult to read. Recently, with the advent of large language models (LLMs), researchers began to explore the potential of LLMs to refine decompiler output. Nevertheless, our study of these approaches reveals their problems, such as introducing new errors and relying on unreliable accuracy validation.
@@
In this paper, we present D-LiFT, an enhanced decompiler-LLM pipeline with a fine-tuned LLM using code quality-aware reinforcement learning. Unlike prior work that overlooks preserving accuracy, D-LiFT adheres to a key principle for enhancing the quality of decompiled code: preserving accuracy while improving readability. Central to D-LiFT, we propose D-Score, an integrated code quality assessment system to score the decompiled source code from multiple aspects, and use it to guide reinforcement learning fine-tuning and to select the best output during inference. In line with our principle, D-Score assigns low scores to any inaccurate output and only awards higher scores for readability to code that passes the accuracy check. Our implementation, based on Ghidra and a range of LLMs, demonstrates significant improvements for the accurate decompiled code from the coreutils and util-linux projects. Compared to baseline LLMs without D-Score-driven fine-tuning, our trained LLMs produce 55.3% more improved decompiled functions, as measured by D-Score. Overall, D-LiFT improves the quality of 68.2% of all the functions produced by the native decompiler.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.240977 www.ndss-symposium.org
I Introduction
@@
Binary decompilation is the process of generating high-level human-comprehensible source code from a binary program. As a fundamental component in security, decompilation supports extensive tasks, such as software reverse engineering [61], vulnerability discovery [40, 22], program patching and hardening [49], and cyber forensics [9]. However, source code generated by decompilers often has problems with syntactic and semantic correctness [37, 75, 11, 16]: A piece of decompiled code may not compile or preserve the semantics of the original binary. Moreover, despite advances in the past decade [7, 72, 12, 35, 74, 69, 68, 8, 33], even with syntactic and semantic correctness, the decompiled code may still be hard for human users to read.
@@
In recent years, with the advent of Large Language Models (LLMs), researchers have started to explore the potential of LLMs to generate and improve decompiled code [66, 56, 25, 29, 18, 28, 64, 67]. For example, to improve decompiled output readability, LLM4Decompile [56] trains new LLMs via supervised fine-tuning (SFT), and DeGPT [25] introduces a three-role mechanism to help verify LLMs’ output during inference. These LLM-assisted decompilation improvement tools have rapidly gained attention. For instance, LLM4Decompile was ranked 4th in GitHub Day Trending.
Original Decompiled Code:
long FUN_0011c6ec(undefined4 param_1,undefined8 param_2,long param_3)
long lVar1;
int *piVar2;
long lStack_30;
lStack_30 = param_3;
while( true ) {
do {
lVar1 = write(param_1,
param_2,lStack_30);
if (-1 < lVar1) {
return lVar1;
piVar2 = (int *)__errno_location();
} while (*piVar2 == 4);
piVar2 = (int *)__errno_location();
if (*piVar2 != 0x16) break;
if (lStack_30 < 0x7ff00001) {
return lVar1;
lStack_30 = 0x7ff00000;
return lVar1;
GPT-o4-mini Improved Code:
#include <unistd.h>
#include <errno.h>
ssize_t write_with_retry(int fd, const void *buf, size_t count) {
const size_t MAX_CHUNK = 0x7FF00000;
size_t to_write = count;
ssize_t result;
while (1) {
result = write(fd, buf, to_write);
if (result >= 0) {
return result;
if (errno == EINTR) {
continue;
if (errno == EBUSY) {
if (to_write > MAX_CHUNK) {
to_write = MAX_CHUNK;
continue;
return result;
Source Code:
ptrdiff_t
safe_rw (int fd, bufptr buf, idx_t count)
for (;;)
ssize_t result = rw (fd, buf, count <= (size_t) -1 ? count : (size_t) -1);
if (0 <= result)
return result;
else if (IS_EINTR (errno))
continue;
else if (errno == EINVAL && SYS_BUFSIZE_MAX < count)
count = SYS_BUFSIZE_MAX;
else
return result;
Figure 1: The code snippet shows the original decompiled code, its corresponding source code, and the version improved by the GPT-o4-mini, where the line 17 of the GPT-o4-mini-refined code is misinterpreted, causing the function’s semantics to diverge from both the original decompilation (line 18) and the true source (line 12). To generate the improved output, we fed GPT-4o-mini (text format, default medium reasoning) the prompt: "help me make the following code more readable without comments:{original decompiled code}".
Unfortunately, applying LLMs to improve decompiled code may trade accuracy for readability. For instance, as shown in Figure˜1, while the GPT-o4-mini [44] improves readability by recovering better variable names and collapsing nested loops into a single loop, it erroneously replaces the condition hex value 0x16 (line 18 of Ghidra’s decompilation) with EBUSY (16 in decimal) instead of the correct constant (EINVAL), introducing a semantic discrepancy. Furthermore, our analysis of the above state-of-the-art LLM-based decompiled code improvement methods also shows significant limitations in preserving the original semantics. For LLM4Decompile, our study (detailed in Section˜VI-A3) reveals that, even when the original decompiled code is error-free, new inaccuracies are found in 93.2% of the functions generated by its fine-tuned LLM. For DeGPT, though it provides a semantic check during inference, its check is non-deterministic and never part of training; it cannot enhance the LLM’s inherent capability to generate more accurate decompiled code. These findings demonstrate a clear need for a more rigorous, LLM-driven decompiled-code enhancement pipeline, one that incorporates a novel training paradigm to produce sturdier models and a rigorous inference-time accuracy check to ensure high-readability and semantics-preserving outputs.
To address these problems, we propose D-LiFT  [^1] D-LiFT” reflects the decompilation pipeline with a “Decompiler” front-end and an “LLM with Fine Tuning” back-end., an enhanced decompiler-LLM pipeline with a fine-tuned LLM using code quality-aware reinforcement learning. D-LiFT takes a baseline LLM, a decompiler, and a set of real-world binaries as inputs, trains a quality-aware LLM as the decompiler’s backend. The enhanced decompiler-LLM pipeline will then output decompiled code of higher quality. Unlike prior work, D-LiFT follows a key principle: preserving accuracy while improving readability, throughout both its training and inference. To this end, D-LiFT first leverages reinforcement learning to fine-tune the baseline LLM and then, at inference, selects the top-scored output from the baseline LLM, the fine-tuned LLM, or the native decompiler.
@@
Central to D-LiFT is D-Score, a novel code quality assessment function, which provides multi-aspect scoring for LLM-generated decompiled code based on accuracy and readability measurements. Specifically, to assess the accuracy of decompiled code, D-Score employs a compiler to generate the syntax feedback and symbolic execution to compare the semantics of the generated code against the corresponding function in the original binary. To assess the readability of decompiled code, D-Score computes a score by applying two established readability metrics to compare the LLM’s output with the original decompiled code. By leveraging D-Score in reinforcement learning, D-LiFT enhances the LLMs’ ability to generate higher-quality decompiled code. Meanwhile, by applying D-Score at inference, D-LiFT reliably selects the best result among the native decompiler, the baseline LLM, and the fine-tuned LLM.
@@
We implement D-LiFT using Ghidra [3] and a number of LLMs, and evaluate D-LiFT using functions from the coreutils [13] and util-linux [59] projects. As measured by D-Score, models fine-tuned by D-LiFT improve the quality of 55.3% more functions compared to the baseline LLMs. Interestingly, we observe a significant “improve-ability gap” between the accurate and inaccurate decompiled functions originally produced by the decompiler. All LLMs, including baseline and fine-tuned models, face challenges in improving the originally inaccurate decompiled code, with improvement rates of just 8.02% for the originally inaccurate functions, in contrast to 86.2% for the originally accurate ones. Overall, for functions that are accurately decompiled, D-LiFT improves the quality (measured by D-Score) of 68.2% of them, where 47.3% of them are improved by the D-LiFT fine-tuned LLM, and only 20.9% by the baseline LLM. We have anonymously released our training scripts and evaluation results at  [^2] https://github.com/XXXXX/xxx/tree/main. In summary, our main contributions include the following:
@@
  • We design D-LiFT, an enhanced decompiler-LLM pipeline with a fine-tuned LLM using code quality-aware RL to improve the quality of the decompiled code, adhering to the principle of preserving accuracy while improving readability.
@@
  • We propose D-Score, an integrated scoring mechanism to quantitatively assess the quality of decompiled code. The framework incorporates existing analytical tools and proven metrics to deliver a multi-aspect evaluation of decompiled code, assessing both syntactic and semantic correctness as well as readability properties to provide effective feedback to the LLM for training and a reliable selection metric for inference.
@@
  • We implement D-LiFT based on Ghidra and three LLMs, and achieve significant decompiled code improvement for commonly used benchmark functions.
II Background and Motivation
II-A Decompiled Code Accuracy
The accuracy of the decompiled code, including syntactic recompilation failures and semantic deviations between the decompiled code and the original source code, was first systematically evaluated using Equivalence Modulo Input (EMI) tests [37] at the function level. Since then, various methods, including symbolic execution [75], random testing [11], fuzzing [70], and manual inspection [16], have been employed to assess decompiled function’s accuracy.
II-B Decompiled Code Readability
Quantitative assessment of code readability has long been a topic in software engineering. Buse and Weimer (B&W) [10] led the way by recruiting 120 human evaluators to rate 100 short code function snippets, then building a mathematical model, featuring metrics like the number of variables per function, to evaluate the readability.
For decompilation, R2I [17] was the first approach to measure readability specifically for decompiled code at the function level. It defines a set of features that can be categorized into five feature groups: code quality, user preference, conflicting features, erroneous syntax, and general features. Analyzing these features extracted from the code from user surveys with 22 participants provides relative readability scores across different decompilers. Nevertheless, since R2I focuses on the comparison across different decompilers, many of its features cannot be generically applied to the LLM-generated code.
II-C LLM application and training
II-C1 Applications of LLMs
@@
Large language models (LLMs) have seen wide adoption and application, particularly in the domain of code-related tasks. Industry has released numerous LLM-powered tools, e.g., GitHub Copilot [19], Google Gemini Code Assistant [20], and Meta LLaMA Coder [51], to streamline different programming tasks. Academia has mirrored this enthusiasm: the number of publications on LLM code programming rose from 11 in 2022 to 75 in 2023 and then to 140 in 2024—a 1,272% increase [27]. In the security domain, people have also leveraged LLMs to improve the decompiled code. For instance, during inference, DeGPT [25] introduces a three-role framework, optimization-scheme referee, rectification-measures advisor, and semantic-check operator, to assist the LLM in producing more readable code. Most recently, people use Model Context Protocol (MCP) servers to infer variable and function names ; one such tool, GhidraMCP [31], garnered 4.3K stars on GitHub in the first month after release.
II-C2 Training LLMs
@@
Instruction tuning uses the supervised learning paradigm, aiming to align the model’s output with a single desired completion. For instance, LLM4Decompile [56] applies it to align the desired output with the original source code.
@@
Reinforcement learning (RL), on the other hand, refines the model through iterative actions and feedback: the model generates candidate outputs (actions), which will be assessed by a reward function that provides feedback, guiding the model toward higher-quality results. In code generation, the feedback usually comes from frameworks, such as unit tests and compilers. Some RL policies, such as PPO [53], also require a single ideal completion to train a critic model that predicts long-term rewards, helping the model choose better actions over time. More recently, Group Relative Policy Optimization (GRPO) [54] achieved promising results on math-related tasks. By normalizing the rewards of candidate outputs instead of relying on a critic model, GRPO eliminates the limitation of accepting only one ideal completion.
II-D Motivations
@@
Although LLM-based decompiled code improvement is increasingly prevalent, it still has several major limitations.
@@
LLMs introduce inaccuracy in the decompiled code. LLMs, operating as probabilistic models that generate text through token-level predictions, systematically introduce inaccuracies in code quality improvement tasks, as evidenced by recent empirical studies documenting pervasive hallucination phenomena [73, 47, 27]. In the decompiled-code improvement scenario, we also observe that LLM-refined output frequently shows new errors. Here, we follow the previous work [37] to define the inaccuracy in the decompiled code as either syntax errors that prevent successful compilation or semantic deviations between the LLM-generated code and the original binary’s behavior. As shown in Table˜III, an average 44.2% of functions that were originally accurate become inaccurate after LLM processing. For the root cause, as case study examples shown later in Section˜VI-D, we observed that LLMs frequently make errors on small details, such as omitting brackets or instructions, or referencing the wrong variable, that are hard to spot yet vital for accuracy.
@@
These observations reveal that though LLMs may improve readability, they often introduce inaccuracy of decompiled code, a factor that prior studies [75, 37, 11, 16, 70] have identified as vital for effective decompilation. Hence, we propose an overarching principle for improving decompiled code quality: preserving accuracy while improving readability.
However, the existing "decompiler-LLM" pipeline does not follow the above principle. Specifically, LLM4Decompile, while effectively boosting LLM’s capability of generating more readable code through training, it compromises the accuracy of its output, as shown in Table˜III. While running its MSSC, a mutated unit test, to check LLM output code semantics among inference, DeGPT uses random inputs that produce non-deterministic semantic check results and omits the check for syntax errors. In light of these limitations, we are motivated to develop a more effective "decompiler-LLM" pipeline, where the LLM backend is specifically designed for improving the readability of the decompiled code while preserving its syntactic and semantic accuracy.
III Design Challenges
Section II-D motivates us to propose a "decompiler-LLM" pipeline without sacrificing accuracy. In this section, we summarize the design-level challenges.